# SPECIAL RELATIVITY

### v9

@@ 3.6 A Clock Resynchronization Procedure
After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration up to velocity $v$. The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation. To get around this difficulty, we observe that simplification is always possible. The trick needed here is to further make a change of the time variable. Suppose that we describe the effect of radiation by working only up to the first order in $v/c$. In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$. After this time shift we obtain the usual wave equation.
Let us now consider the case of an arbitrary velocity along the $x$-axis. We have, then, a general method for finding solution of electrodynamics problem in the case of the absolute time coordinatization. The new independent variables $(x_{L},t_{L})$ can be expressed in terms of the old independent variables $(x,t)$:
@@
Since the change variables completed by the Galilean transformation is mathematically equivalent to the Lorentz transformation, it obviously follows that transforming to new variables $x_{L},t_{L}$ leads to the usual Maxwell’s equations. In particular, the wave equation Eq. (12) transforms into
@@
In the new variables the velocity of light emitted by a moving source is constant in all directions, and equal to the electrodynamics constant $c$.
@@
The overall combination of the Galilean boost and variable changes actually yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame, but in this context this transformation are only to be understood as useful mathematical device, which allow one to solve the electrodynamics problem in the choice of absolute time synchronization with minimal effort.
@@
We can now rise an interesting question: do we need to transform the results of the electrodynamics problem solution into the original variables? We state that the variable changes performed above have no intrinsic meaning - their meaning only being assigned by a convention. In particular, one can see the connection between the time shift $t\to t+xv/c^{2}$ and the issue of clock synchrony. Note that the final change in the scale of time and spatial coordinates is unrecognizable also from a physical viewpoint. It is clear that the convention-independent results of calculations are precisely the same in the new variables. As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original variables.
@@
Consider now two light sources say ”1” and ”2”. Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$. The question now arises how to assign a time coordinate to the lab reference frame. We have a choice between an absolute time coordinate and a Lorentz time coordinate. The most natural choice, from the point of view of connecting to the laboratory reality, is the absolute time synchronization. In this case simultaneity is absolute, and for this we should prepare, for two sources, only one set of synchronized clocks in the lab frame. On the other hand, Maxwell’s equations are not form-invariant under Galilean transformations, that is, their form is different on the lab frame. In fact, the use of the absolute time convention, implies the use of much more complicated field equations, and these equations are different for each source. Now we are in the position to assign Lorentz coordinates. The only possibility to introduce Lorentz coordinates in this situation consists in introducing individual coordinate systems (i.e. individual set of clocks) for each source. It is clear that if operational methods are at hand to fix the coordinates for the first source, the same methods can be used to assign values to the coordinates for the second source and these will be two different Lorentz coordinate systems.
Suppose now that the second secondary source (e.g. mirror or pupil detection) interacts with the light beam radiated from the first source. The peculiarity of this problem with the viewpoint of the special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame.
@@
The utilization of the theory of relativity becomes indispensable when we consider optical phenomena associated with a motion of light sources. So far we have considered the covariant and non covariant ways to solve the problem of radiation by a moving source in an inertial frame. Let us consider the acceleration of a dipole source in the lab inertial frame up to velocity $v$ along the $x$-axis. The question arises how to assign synchronization in the lab frame after the source acceleration. Before acceleration we picked a Lorentz coordinate system. Without changing synchronization in the lab frame after the source acceleration we have a complicated situation as concerns electrodynamics of moving charges. As a result of such boost, the transformation of time and spatial coordinates has no form of a Lorentz transformation. The Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned. In order to keep the Lorentz coordinatiyation after acceleration, one needs to change the space-time coordinate system by introducing new variables $(x_{L},y_{L},z_{L},t_{L})$, Eq. (13). In the new variables the velocity of light emitted by a moving source is equal to the electrodynamics constant $c$ in all directions.
@@
The aberration of light and Doppler effect are practical cases of study for illustration the difference between covariant and non-covariant approaches. The present treatment is designed to try and bring out the role of the Lorentz time transformation. The most convenient way for calculating radiation is to make use of the formulae for Lorentz transition from one inertial frame of reference to another. It is convenient to introduce the reference frame $S^{\prime}$ fixed to the source. Then the problem solved as follows. In the frame $S^{\prime}$ the 4-vector of the light beam is specified, i.e. the frequency and the propagation direction of light are known. The frequency of light in the frame $S$ is easily to find using the Lorentz transformation formulae. Specifically, let the radiation have a dipole character and frequency $\omega_{0}$ in the frame of reference in which the source is at rest. Then in the laboratory frame of reference $S$ in which the source as a whole moves with velocity $v$ and the radiation comes along the velocity direction we observe the so-called radial Doppler effect. It is described by the well-known formula $\omega=\omega_{0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$. The effect of the factor $\sqrt{1-v^{2}/c^{2}}$ can be summarized in the following statement: on the moving object time is flowing slower than expected (time dilation).
@@
It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily leads to mistake. For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’ equations are compatible with the result of conventional particle tracking. This method was incorrectly extended to the case of non-collinear geometry.
@@
At this point a reasonable question arises: why the same method gives incorrect result in the case of radiation of light from a source moving perpendicular to its radiated light beam propagation. It is not difficult to see that the peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e plane of simultaneity). Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
@@
So far we have dealt with the longitudinal effect when the motion takes place along the straight line through the source and the receiver. In the transverse case the motion is perpendicular to this direction and the source is accelerated along the radiation wavefront (i.e. initial plane of phase simultaneity). The orientation of the radiation wavefront is therefore no longer something absolute and depends on the clock synchronization convention. Let us first consider a Lorentz transformation. The explanation consists in using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame. It is clear, that wavefront phases which are simultaneous in $S^{\prime}$, but occur at different $x^{\prime}$ locations, are not simultaneous in $S$. If make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$ and the effect of this transformation is just a rotation of the radiation wavefront in the lab frame on the angle $v/c$ in the first order approximation. According to Maxwell’s electrodynamics, radiation is always emitted in the direction normal to the radiation wavefront. Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light. This remarkable phenomenon is what we shall discuss in the next chapter. According to the conventional (3+1) approach, the simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame. Thus it seems as if the conventional (3+1) approach is unable to account for the geometric phenomenon of aberration. In making use of the Maxwell’s equations and the absolute time transformation we obtain no deflection at all of the wavefront plane and the energy transport direction.
@@
Let us now return to our consideration of the conventional coupling between Maxwell’s equations and particle trajectories in the ”single lab frame”. According to the theory of special relativity, there is no objection to the standard description of the optical phenomena associated with a motion of light source in the case when the radiation comes along the velocity direction. For example, for collinear geometry, the conventional coupling fields and particles may be a useful approach for analysis of the Doppler effect.
In order to understand the relativistic red shifting of the light source in a moving system, we have to watch the machinery of the source and see what happens when it is moving. Since that is rather difficult, we shall take a very simple a kind of source, but it will work in principle.
@@
According to covariant approach, the various relativistic kinematics effects concerning to the dipole radiation setup, turn up in successive orders of approximation. In the first order $(v/c)$. - relativity of simultaneity. In the second order $(v/c)^{2}$. - time dilation. The first order kinematics term $(v/c)$ plays an essential role only in the description of the dipole radiation in the perpendicular geometry. In the case of collinear geometry, a motion of the dipole source, according to the theory of relativity, influences the kinematics terms of the second order $(v/c)^{2}$ only. The ignorance of this distinction between collinear and non collinear boosts is the source of much confusion in the literature.
@@
In the non-covariant approach, a solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. This means that, for instance, within the lab frame the motion of particles looks precisely the same as predicted by classical mechanics, with its absolute time. The relativity of simultaneously (i.e. mixture of positions and time) do not have a place in this description. Such (the absolute time) method is suitable to account for the outcome of the experiments in collinear geometry. We argue that this algorithm for solving usual Maxwell’s equations in the lab frame is not applicable in the transverse case. However, the Maxwell’s equations in the lab frame are compatible only with covariant trajectories calculated by using Lorentz coordinates, therefore including such relativistic features as relativity of simultaneity.
3.8 Absolute Time Synchronization in a Laboratory Frame
@@
The utilization of the electrodynamics in the absolute time coordinatization becomes indispensable when we use non-covariant (3+1) dimensional approach (i.e. old kinematics) to relativistic particle dynamics. According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab time $t$ as independent variable. The relativistic kinematics effects do not have a place in this description. Such approach to relativistic particle dynamics is actually based on the use of the absolute time synchronization assumption in the lab frame. If one wants to use the corrected Newton equations (in other words, using time as independent variable), only solution of the electrodynamics equations in non-covariant form (i.e. in the absolute time coordinatization) gives the correct coupling between electrodynamics equations and particle trajectories in the ”single frame”. That is the reason why the absolute time synchronization convention is extremely important from the applied point of view.
In the Chapter 16 we shall discuss one of the consequences of our aberration of light theory, the radiation from an X-ray Free Electron Laser. Using XFEL example we can demonstrate that the coupling fields and particles in the conventional approach is intrinsically incorrect. The usual XFEL theory based on the use of old Newtonian kinematics for particle dynamics and the Einstein’s kinematics for the electrodynamics. In fact, the usual theoretical treatment of relativistic particle dynamics involves only a corrected Newton’s second law and is based on the use Galilean edition of velocities. This coupling of Maxwell’s equations and corrected Newton’s equation is commonly accepted as useful method in accelerator physics and, in particular, in numerical calculations of radiation properties. For rectilinear motion of the (ultra-relativistic) modulated electron beam, non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking.
Let us now move on to consider the predictions of the existing XFEL theory in the case of non-collinear electron beam motion. As well-known result of conventional particle tracking states that after an electron beam is kicked by a weak dipole magnet there is a change in the trajectory of the electron beam, while the orientation of the modulated wavefront remains as before. In the absolute time coordinatization the simultaneity of ivents have absolute character. The absolute character of temporal coincidence of events is a consequence of the (hidden) absolute time synchronization convention. According to this old kinematics, the modulation wavefront remains unvaried. In other words, the kick results in a difference between the directions of the electron motion and the normal to the modulation wavefront (i.e. the kick results in a wavefront tilt).
In existing XFEL theory the wavefront tilt is considered as real. However, in the ultra-relativistic limit, the wavefront tilt has no exact objective meaning at all since, due to finiteness of the speed of light, we cannot specify any experimental method by which this tilt could be ascertained. According to conventional (3+1) space-time treatment, a transverse kick does not change the orientation of a modulation wavefront, and hence suppresses the radiation emitted in the direction of the electron beam motion. In fact, according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront. This is because Maxwell’s theory has no intrinsic anisotropy. The electrons lying on the plane of simultaneity (i.e. on the modulation wavefront) gives rise to spherical radiation wavelets, and these combine according to Huygens’ principle to form what what is effective a radiated wave. If the plane of simultaneyty is $xy$-plane (i.e. beam modulation wavefront is perpendicular to the $z$-axis), then the Huygens’s construction shows that plane radiation wavefront will be emitted along $z$-direction.
Only solution of the electrodynamics equations in non-covariant form gives correct coupling between the usual particle trajectories and the fields in the lab frame. The non-covariant approach within the framework of both mechanics and electrodynamics predicts an effect in complete contrast to the conventional treatment.
In the Chapter 16 we demonstrate that our Galilean transformed electrodynamics within the lab frame (see Eq.11) says that by making measurement on the coherent radiation, one can observe only radiation in the kicked direction. Experiments (involving specially kicked modulated electron beam) show that this prediction is, in fact, true. The results of XFEL experiments demonstrated that even the direction of emission of coherent undulator radiation is beyond the predictive power of the conventional theory. A starting point in our study of relativity was this complete disagreement between XFEL theory and experiment.
At this point, a reasonable question arises: since XFELs are designed without accounting for the special relativity, how can they actually operate? All existing XFEL codes are based on a model in which the modulated electron beam moves only along the undulator axis. In a typical configuration for an XFEL, the orbit of a modulated electron beam is controlled to avoid large excursions from the undulator axis. However, random errors in the focusing system can cause angular trajectory errors (or ”kicks”). Analysis of the trajectory errors on the XFEL amplification process showed that any XFEL undulator magnetic field must satisfy stringent requirements. However, studies of this critical aspect in the design of a XFEL sources are based on an incorrect coupling of fields and particles. The pleasant surprise is that the tolerances predicted are more stringent than they need be according to our corrected XFEL theory. This can be considered as one of the reason for the exceptional progress in XFEL developments over last decades.
It is generally assumed that the usual Maxwell’s equations and corrected Newton’s second law can explain all experiments that are performed in a single inertial (lab) frame. According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. There are many physicists who have already received knowledge about special relativity from textbooks and who would say, ”For those who want to learn just enough about special relativity so they can solve problems - it just changes Newton’s laws by introducing a correction factor to the mass.” This is misconception. We cannot take old kinematics for mechanics and Einstein’s kinematics for electrodynamics.
3.10 Bibliography and Notes
1. According to the thesis of conventionality of simultaneity [13, 14, 15, 16, 17], simultaneity of distant events is a conventional matter, as it can be legitimately fixed in different manners in any given inertial reference frame. To quote e.g. Moeller [13]: ”All methods for the regulation of clocks meet with the same fundamental difficulty. The concept of simultaneity between two events in different places obviously has no exact objective meaning at all, since we cannot specify any experimental method by which this simultaneity could be ascertained. The same is therefore true also for concept of velocity.” Let us illustrate a particular special example of the simultaneity convention. Let the synchronization of clocks in different spatial points be provided by light signals having, respectively, velocity $c_{1}$ in the direction parallel to the positive axis $x$, and velocity $c_{2}$ in the opposite direction. Then, a light signal sent from point $A$ at time $t_{A}$ will arrive at point $B$ at time $t_{B}=t_{A}+x_{AB}/c_{1}$. The reflected signal will arrive back at point $A$ at time $t^{\prime}_{A}=t_{B}+x_{AB}/c_{2}$. From these two expressions we get $t^{\prime}_{A}-t_{A}=x_{AB}(c_{2}+c_{1})/(c_{1}c_{2})$. Summing up we have $t_{B}=t_{A}+c_{2}(t^{\prime}_{A}-t_{A})/(c_{1}+c_{2})$. So we come to the synchronization first proposed in [14]: $t_{B}=t_{A}+\epsilon(t^{\prime}_{A}-t_{A})$, where $\epsilon=c_{2}/(c_{2}-c_{1})$. After substituting in this expression Einstein choice $c_{1}=-c_{2}=c$ we get $\epsilon=1/2$. This time order, as fixed by the standard synchronization, is frame dependent. This is the well-known thesis of relativity of simultaneity. The conventional nature of distant simultaneity in special relativity is not to be confused with the relativity of simultaneity. Clearly, the conventionality of simultaneity within a single inertial frame is quite distinct from the relativity of simultaneity in Einstein’s synchronization. If $\epsilon\neq 1/2$, then speed of light from $A$ to $B$ differs from the speed of light from $B$ to $A$. Although the Einstein synchronization choice, is preferred by physicists, it is nothing more ”physical” than any other. A nice pedagogical (but artificial) example of a non-standard synchronization is clock synchronization [11] in which $\epsilon=0$. To any convention on the simultaneity there will correspond a definite choice of the coordinate system in an inertial coordinate system of reference of the Minkowski space-time. A particular very unusual choice of coordinates, the absolute time coordinate choice, is exploited in this book. In our case, when time is absolute, we have $c_{1}=c+v$, $c_{2}=-c+v$. Substituting into the expression for $\epsilon$ we obtain $\epsilon=(1/2)(1-v/c)$.
2. In the text first published in 1923, Eddington discussed, apparently for the first time, a procedure for synchronization using slow transport of clocks [18]. The details can be found in review [8] ( see also [19]).
3. We already pointed that we have empirical access only to the round-trip average speed of light. An empirical test of any property of the one-way speed of light is not possible. Many authors of textbooks still attribute a reality status of the one-way speed of light. To quote Hrasko [20]: ”It is sometimes claimed that Einstein synchronization of distant clocks $A$ ans $B$ is circular. The argument is very simple: Einstein synchronization is based on the equality of light velocity on the path from $A$ to $B$ and back from $B$ to $A$, but measurement of light velocity in one direction between to distant points is impossible unless the clocks at these points have already been synchronized. This argument is, however, fallacious. It is true that one-way measurement of light velocity can be performed only if clocks at the endpoints are synchronized correctly. But since they need not show the correct coordinate time, they can be synchronized without light signals by transporting them from common site in a symmetrical manner. The procedure consists of following steps: … . As we see, the thought experiment described is capable to prove constancy of light speed if it is true, or to disprove it if it is false. It provided, therefore, solid logical foundation for Einstein’s synchronization prescription.” This logical argument is incorrect. Slow clock transport synchronization is equivalent to Einstein’s synchronization in inertial system where the light source is at rest.
4. For a general discussion of correspondence between passive and active transformations we suggest reading the paper [21].
@@ 3.10 Bibliography and Notes
4.1 Introductory Remarks
The effect of light aberration in an inertial frame of reference is usually understood as a change in the direction of light propagation ascribed to boosted light sources. The phenomenon of aberration of light in an inertial frame of reference by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the textbooks. Thus, in order to keep the mathematical complexity of the discussion to a minimum, we will describe the effect of aberration by working only up to the first order. The appearance of relativistic effects in optical phenomena does not depend on a large speed of the radiation sources. Light is always a relativistic object, no matter how small the ratio $v/c$ may be.
The subject of this chapter is the physical influence of the optical instrument on the aberration of light measurement. First, we will demonstrate that when one has finite-aperture mirror moving transversely and the plane wave of light is falling normally on the mirror, there is the aberration (deviation of the energy transport) for light reflected from the mirror. The next problem to be discussed in this chapter is of more practical importance. According to the Babinet’s principle, the problem of reflection from a transversely moving (finite-aperture) mirror complementary to the problem of transmission through a hole in the moving opaque screen. One case of rather great interest is that which corresponds to transmission through a moving end of the telescope barrel. It is shown that the (electromagnetic) wave theory approach to the aberration of light in an inertial frame leads to a substantially different result for energy transport than the predicted by conventional theory.

### v10

@@ 3.5 A Clock Resynchronization Procedure
After the Galilean boost in the case of the absolute time coordinatization we can see that the homogeneous wave equation for the field in the lab frame has nearly but not quite the usual, standard form that takes when there is no acceleration. The main difference consists in the crossed term $\partial^{2}/\partial t\partial x$, which complicates the solution of the equation. To get around this difficulty, we observe that simplification is always possible. The trick needed here is to further make a change of the time variable. Suppose that we describe the effect of radiation by working only up to the first order in $v/c$. In this approximation, it is always possible to introduce a time transformation $t\to t+xv/c^{2}$. After this time shift we obtain the usual wave equation  [^14] Rotation of the $x$-axis without changing the orientation of the $t$-axis is described by a Galilean transformation $x^{\prime}=x-vt$, $t^{\prime}=t$. Rotation of the $t$-axis without changing the orientation of the $x$-axis is described by the transformation $x^{\prime}=x$, $t^{\prime}=t+xv/c^{2}$. In the first-order approximation, this transformation mathematically is similar to the the Galilean transformation $x^{\prime}=x+vt$, $t^{\prime}=t$.
Let us now consider the case of an arbitrary velocity along the $x$-axis. We have, then, a general method for finding a solution of the electrodynamics problem in the case of absolute time coordinatization. The new independent variables $(x_{L},t_{L})$ can be expressed in terms of the old independent variables $(x,t)$:
@@
Since the change variables completed by the Galilean transformation is mathematically equivalent to the Lorentz transformation, it follows that transforming to new variables $x_{L},t_{L}$ leads to the usual Maxwell’s equations. In particular, the wave equation Eq. (12) transforms into
@@
In the new variables, the velocity of light emitted by a moving source is constant in all directions, and equal to the electrodynamics constant $c$.
@@
The overall combination of the Galilean boost and variable changes yields the Lorentz transformation in the case of absolute time coordinatization in the lab frame, but in this context, this transformation is only to be understood as a useful mathematical device, which allows one to solve the electrodynamics problem in the choice of absolute time synchronization with minimal effort.
@@
We can now raise an interesting question: do we need to transform the results of the electrodynamics problem solution into the original variables? We state that the variable changes performed above have no intrinsic meaning - their meaning only being assigned by a convention. In particular, one can see the connection between the time shift $t\to t+xv/c^{2}$ and the issue of clock synchrony. Note that the final change in the scale of time and spatial coordinates is unrecognizable also from a physical viewpoint. It is clear that the convention-independent results of calculations are precisely the same in the new variables. As a consequence, we should not care to transform the results of the electrodynamics problem solution into the original variables.
@@
Consider now two light sources say ”1” and ”2”. Suppose that in the lab frame the velocities of ”1” and ”2” are $\vec{v}_{1}$, $\vec{v}_{2}$ and $\vec{v}_{1}\neq\vec{v}_{2}$. The question now arises of how to assign a time coordinate to the lab reference frame. We have a choice between an absolute time coordinate and a Lorentz time coordinate. The most natural choice, from the point of view of connecting to the laboratory reality, is absolute time synchronization. In this case, simultaneity is absolute, and for this, we should prepare, for two sources, only one set of synchronized clocks in the lab frame. On the other hand, Maxwell’s equations are not form-invariant under Galilean transformations, that is, their form is different in the lab frame. The use of the absolute time convention implies the use of much more complicated field equations, and these equations are different for each source. Now we are in the position to assign Lorentz coordinates. The only possibility to introduce Lorentz coordinates in this situation consists of introducing individual coordinate systems (i.e. individual sets of clocks) for each source.
Suppose now that the second secondary source, for instance, a tangentially moving mirror, interacts with the light beam radiated from the first source at rest. The peculiarity of this problem with the viewpoint of special relativity is that here we cannot prepare for sources a common Lorentz coordinatization in the single inertial frame  [^15] It is clear that the electrodynamics equations must be identical for incoming and reflected waves. The time under Einstein’s synchronization for the moving mirror is readily obtained by introducing the time shift $xv/c^{2}$ in the first-order approximation. In this synchronization, we describe the reflection using Maxwell’s equations. This time shift has the effect of rotation of the radiation wavefront on the angle $v/c$ for both sources. Let us discuss the electrodynamics of the source at rest. In the absolute time coordinatization, the observer describes the light beam from the emitter at rest using the diagonal wave equation. The new crossed term that has to be put into the wave equation due to time shift leads to the non-diagonal wave equation (see footnote in Section 3.5).(see the next section for more details).
@@
So far we have considered the covariant and non covariant ways to solve the problem of radiation by a moving source in an inertial frame. Let us consider the acceleration of a dipole source in the lab inertial frame up to velocity $v$ along the $x$-axis. The question arises of how to assign synchronization in the lab frame after the source acceleration. Before acceleration, we picked a Lorentz coordinate system. Without changing synchronization in the lab frame after the source acceleration we have a complicated situation concerning the electrodynamics of moving charges. As a result of such a boost, the transformation of time and spatial coordinates has no form of a Lorentz transformation. Maxwell’s equations can be applied in the lab inertial frame only in the case when Lorentz coordinates are assigned. To keep the Lorentz coordinatization after acceleration, one needs to change the space-time coordinate system by introducing new variables $(x_{L},y_{L},z_{L},t_{L})$, Eq. (13). In the new variables, the velocity of light emitted by a moving source is equal to the electrodynamics constant $c$ in all directions.
@@
The aberration of light and the Doppler effect are practical cases of study to illustrate the difference between covariant and non-covariant approaches. The most convenient way for calculating radiation is to make use of the formulae for Lorentz transition from one inertial frame of reference to another. Specifically, let the radiation have a dipole character and frequency $\omega_{0}$ in the frame of reference $S^{\prime}$ in which the source is at rest. Then in the laboratory frame of reference $S$ in which the source as a whole moves with velocity $v$ and the radiation comes along the velocity direction we observe the so-called radial Doppler effect. It is described by the well-known formula $\omega=\omega_{0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$. The effect of the factor $\sqrt{1-v^{2}/c^{2}}$ can be summarized in the following statement: on the moving object, time is flowing slower than expected (time dilation).
@@
It is interesting to note that, the use of the conventional coupling of Maxwell’s equations and corrected Newton’s equation for the calculation of the radiation from a moving source does not necessarily lead to a mistake. For rectilinear motion of the source and the emitted light beam, the non-covariant and covariant approaches produce the same trajectories, and Maxwell’s equations are compatible with the result of conventional particle tracking. This method was incorrectly extended to the case of non-collinear geometry.
@@
At this point, a reasonable question arises: why the same method gives incorrect results in the case of radiation of light from a source moving perpendicular to its radiated light beam propagation? It is not difficult to see that the peculiarity of the collinear geometry where there is a source moving along the same line as the radiated beam is that here the velocity is perpendicular to the plane of radiation wavefront (i.e. plane of simultaneity). Thus, for collinear motion, the plane of simultaneity in the absolute time coordinatization will have the same orientation for the Lorentz coordinatization.
@@
In the transverse case, the motion is perpendicular to this direction, and the source is accelerated along the radiation wavefront (i.e. initial plane of phase simultaneity). The orientation of the radiation wavefront is therefore no longer something absolute and depends on the clock synchronization convention. Let us first consider a Lorentz transformation. The explanation consists of using a Lorentz boost to describe the uniform translation motion of the light source in the lab frame. It is clear, that wavefront phases which are simultaneous in $S^{\prime}$ but occur at different $x^{\prime}$ locations, are not simultaneous in $S$. If make a Lorentz boost, we automatically introduce a time transformation $t^{\prime}=t-xv/c^{2}$, and the effect of this transformation is just a rotation of the radiation wavefront in the lab frame on the angle $v/c$ in the first-order approximation. According to Maxwell’s electrodynamics, radiation is always emitted in the direction normal to the radiation wavefront. Then, the radiated light beam is propagated at the angle $v/c$, yielding the phenomenon of the aberration of light. This remarkable phenomenon is what we shall discuss in the next section. According to the conventional (3+1) approach, the simultaneity is absolute and there is no mixture of positions and time when sources change their velocities in the inertial frame. Thus it seems as if the conventional (3+1) approach is unable to account for the geometric phenomenon of aberration. In making use of Maxwell’s equations and the absolute time transformation we obtain no deflection at all of the wavefront plane and the energy transport direction.
@@
Let us now return to our consideration of the conventional coupling between Maxwell’s equations and particle trajectories in the ”single lab frame”. According to the theory of special relativity, there is no objection to the standard description of the optical phenomena associated with a motion of a light source in the case when the radiation comes along the velocity direction. For example, for collinear geometry, the conventional coupling fields and particles may be a useful approach for the analysis of the Doppler effect.
@@
The relativistic motion of an electron is described by $d\vec{p}/dt=e\vec{v}\times\vec{H}/c$ with $\vec{p}=m\vec{v}/\sqrt{1-v^{2}/c^{2}}$. This equation can be written as $d\vec{v}/dt=\vec{v}\times\vec{\omega}_{L}$ where $\vec{\omega}_{L}=e\vec{H}\sqrt{1-v^{2}/c^{2}}/(mc)=\vec{\omega}_{L0}\sqrt{1-v^{2}/c^{2}}$ is the relativistic Larmor frequency. When the source is accelerated, the speed of electrons is increased, and therefore the mass is also increased and the electron is heavier. The emissivity presents a spectrum in which the frequency is given by $\omega=\omega_{L0}\sqrt{1-v^{2}/c^{2}}/(1-v/c)$. The frequency is multiplied by the relativistic Larmor frequency accounting at the same time through the denominator for the Doppler effect caused by the motion parallel to the magnetic field.
We can therefore conclude that the frequency of oscillations of an electron is slowed down if the source is made to move with a velocity $v$. At first sight, the machinery of the source described does not involve the effect of relativistic time dilation at all. But it is there in the assumption that the mass of a moving object is equal to its relativistic mass $m_{r}=m/\sqrt{1-v^{2}/c^{2}}$.
@@
According to the covariant approach, the various relativistic kinematics effects concerning to the dipole radiation setup, turn up in successive orders of approximation. In the first order $(v/c)$. - relativity of simultaneity. In the second order $(v/c)^{2}$. - time dilation. The first-order kinematics term $(v/c)$ plays an essential role only in the description of the dipole radiation in the perpendicular geometry. In the case of collinear geometry, a motion of the dipole source, according to the theory of relativity, influences the kinematics terms of the second order $(v/c)^{2}$ only.
@@
In the non-covariant approach, a solution of the dynamics problem in the lab frame makes no reference to Lorentz transformations. This means that, for instance, within the lab frame the motion of particles looks precisely the same as predicted by classical mechanics, with its absolute time. Such (the absolute time) method is suitable to account for the outcome of the experiments in collinear geometry. We argue that this algorithm for solving Maxwell’s equations in the lab frame is not applicable in the transverse case.
@@
The utilization of the electrodynamics in absolute time coordinatization becomes indispensable when we use a non-covariant (3+1) dimensional approach (i.e. old kinematics) to relativistic particle dynamics. According to conventional particle tracking, the dynamical evolution in the lab frame is based on the use of the lab time $t$ as the independent variable. The relativistic kinematics effects do not have a place in this description. Such an approach to relativistic particle dynamics is based on the use of the absolute time synchronization assumption in the lab frame. If one wants to use the corrected Newton equations (in other words, using time as the independent variable), only the solution of the electrodynamics equations in non-covariant form (i.e. in the absolute time coordinatization) gives the correct coupling between electrodynamics equations and particle trajectories in the ”single frame”.
@@
The effect of light aberration in an inertial frame of reference is usually understood as a change in the direction of light propagation ascribed to boosted light sources. The phenomenon of aberration of light in an inertial frame of reference is by no means simple to describe, even in the first order in $v/c$: a large number of incorrect results can be found in the textbooks. Thus, to keep the mathematical complexity of the discussion to a minimum, we will describe the effect of aberration by working only up to the first order. The subject of this section is the physical influence of the optical instrument on the aberration of light measurement. First, we will demonstrate that when one has a finite-aperture mirror moving parallel to itself and the plane wave of light falls normally on the mirror, there is an aberration (deviation of the energy transport) for light reflected from the mirror. The next problem to be discussed in this section is of more practical importance. According to Babinet’s principle, the problem of reflection from a tangentially moving (finite-aperture) mirror is complementary to the problem of transmission through a hole in the moving opaque screen. One case of rather great interest is that which corresponds to transmission through a moving end of the telescope barrel.